A shipping container is 24 feet long, 5 feet wide and 6 feet high. (a) calculate its volume in cubic yards. (b) express this volume in cubic meters.
1
Expert's answer
2011-02-17T05:47:13-0500
Let's fing the volume in cubic feets.
V = 24*5*6 = 720 feet3.
In one cubic foot there is 0.037 cubic yards. So
V[in yards3] = 720*0.037 = 26.64 yard3
In one cubic foot there is 0.0283 m3:
V[in m3] = 720 *0.0283 = 20.376 m3.